OpenStudy (anonymous):

How do i calculate the number of neutrons in uranium when i have 92 protons and a specific charge of 3.7*10^7?

OpenStudy (anonymous):

no of protons = no of electron and no of neutrons = no of electrons + no of protons.

OpenStudy (anonymous):

so we know the specific charge = (3.7x10^7) so we know the charge is 92(1.6 X10^-19)/ mass which is (92+x)(1.67 x10^-27) so we put mass on oneside and numericals on the other to get: (92 + X) (1.67x10^-27)=3.97x10^-25 3.97..... is (92(1.6x10^-19)) / (3.7x10^7) divide it out 92 + x = 238 ish x = 238-92 X= 146 http://www.thestudentroom.co.uk/showthread.php?t=1282309 (jclark_76)

OpenStudy (unklerhaukus):

units of specific charge = charge / unit mass

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

Where do i get x from?

OpenStudy (anonymous):

x stands for the number of neutrons

OpenStudy (anonymous):

Ok so i have done the calculations and i have 3.97*10^-25. How do i calculate the 238 for the number of neutrons?

OpenStudy (unklerhaukus):

what are the units of that number

OpenStudy (anonymous):

3.97*10^-25 = (92(1.6x10^-19)) / (3.7x10^7)

OpenStudy (anonymous):

How do you come to the answer of 238?

OpenStudy (anonymous):

Can you explain how we come to 238 brent0423?

OpenStudy (anonymous):

Yes i know that but how do you go from this calculation 3.97*10^-25 = (92(1.6x10^-19)) / (3.7x10^7) to X = 238 ?

OpenStudy (anonymous):

How do we arrive at x= 238? When all i have is3.97*10^-25 ?

OpenStudy (anonymous):

Am i missing a calculation or is there more to add?

OpenStudy (unklerhaukus):

\[x\neq 238\] it is very hard to see what you are doing when you don't have units floating around

OpenStudy (anonymous):

Its ok posting threads to different sites for the answers but if you cannot explain the answer of how we arrive at 238 from this 3.97*10^-25 = (92(1.6x10^-19)) / (3.7x10^7) to X = 238 i have given all the units!

OpenStudy (anonymous):

(1.67 x10^-27) so we put mass on oneside and numericals on the other to get: (92 + X) (1.67x10^-27)=3.97x10^-25 3.97..... is (92(1.6x10^-19)) / (3.7x10^7) divide it out 92 + x = 238 ish Where do i get X = 238 ?

OpenStudy (anonymous):

wait one minute and i will provide a step by step solution

OpenStudy (anonymous):

Ok thankyou

OpenStudy (anonymous):

no worries, im sorry that the link i provided was misleading

OpenStudy (anonymous):

Its ok i just need a better explanation

OpenStudy (anonymous):

protons plus neutrons equals 238 so if you have 92 protons you would have to have 146 neutrons

OpenStudy (anonymous):

Yes i know but there are two types of uranium 235 & 238 so how would i know which one to choose?